OpenStudy (horsegal244):

James left a bin outside in his garden to collect rainwater. He notices that 1 over 4 gallon of water fills 2 over 3 of the bin. Write and solve an equation to find the amount of water that will fill the entire bin. Show your work. Explain your answer in words.

OpenStudy (horsegal244):

@Michele_Laino

OpenStudy (michele_laino):

in order to fill completely one bin, we need \[\frac{3}{2} \times \frac{1}{4}\] gallon

OpenStudy (horsegal244):

ok

OpenStudy (michele_laino):

of rainwater

OpenStudy (michele_laino):

so the requested amount Q of water is : \[Q = \frac{3}{2} \times \frac{1}{4} = ...?\] please complete the computation above

OpenStudy (horsegal244):

12/2

OpenStudy (michele_laino):

hint: